OpenStudy (nurali):

In a roll of a die, let a "six" be the face with six spots. Consider the two events below. Event A: more than 16,000 sixes in 100,000 rolls Event B: more than 160,000 sixes in 1,000,000 rolls. Pick the right option. P(A) = P(B) P(A) > P(B) P(A) < P(B)

OpenStudy (perl):

first part Let X = # of 6's in 100,000 rolls X is a binomial random variable , n=100 000 , p = 1/6 P(A) = P ( X > 16 000) second part Let Y = # of 6's in 1,000,000 rolls n= 1 000 000 , p = 1/6 P(B) = P( Y > 160 000 )

OpenStudy (perl):

im not supposed to give the answer :o

OpenStudy (perl):

P(A) = P ( X > 16 000) = 1 - binomcdf( 100 000 , 1/6, 16 000 )

OpenStudy (perl):

ok, its easier if we simplify this problem

OpenStudy (perl):

P(A) < P(B)

OpenStudy (nurali):

Thanks it is correct.